# On the Inductive Biases of Demographic Parity-based Fair Learning Algorithms

### v1

@@ 5 A Distributionally Robust Optimization Approach to DP-based Fair Learning
$\min_{f\in\mathcal{F}}\max_{Q_{S}\mathrel{\mathop{:}}d(Q_{S},P_{S})\leq\delta}\mathbb{E}_{P_{\mathbf{X},Y|S}\cdot Q_{S}}\Bigl{[}\ell_{0/1}\bigl{(}\widehat{Y},Y\bigr{)}\Bigr{]}+\lambda\rho\bigl{(}\widehat{Y},S\bigr{)}$ (5)
According to this formulation, we solve the Lagrangian version of optimization problem (1) when $S$’s marginal distribution $q_{S}$ leads to the worst-case fair-regularized risk function in a distribution ball $\bigl{\{}q_{S}\mathrel{\mathop{:}}\,d(q_{S},p_{S})\leq\delta\bigr{\}}$ where $d$ is a distance measure between probability distributions. In this formulation, we consider assigning different weights to samples with different sensitive attributes, which may result in different majority sensitive attributes. Since we are optimizing the worst-case performance over the distribution ball with a $\delta$ radius, the inductive biases discussed in the previous would become less effective under a greater $\delta$.
The proposed SA-DRO formulation results in Algorithm 1 which applies projected gradient descent ascent (GDA) to solve the minimax optimization problem in (5). Here, we use a parameterized classifier $f_{\mathbf{w}}$ to apply a gradient-based training algorithm. Also, the distance $d$ can be chosen as any standard $f$-divergence. In our experiments, we attempted the $\chi^{2}$-divergence divergence, which has been well-explored in the literature [24, 25, 26].
@@
1:  Input: Training data $\{(\mathbf{x}_{i},y_{i},s_{i})_{i=1}^{n}\}$, parameters $\lambda,\delta\geq 0$, divergence $d$, dependence measure $\rho$, stepsizes $\alpha_{w},\alpha_{q}>0$
@@
${g_{\mathbf{q}}}_{s}=\frac{1}{n}\sum_{i\mathrel{\mathop{:}}s_{i}=s}\bigl{[}\ell\bigl{(}f_{\mathbf{w}}(\mathbf{x}_{i}),y_{i}\bigr{)}\bigr{]}+\lambda\frac{\partial\rho\bigl{(}f_{\mathbf{w}}(\mathbf{x}_{1\mathrel{\mathop{:}}n}),s_{1\mathrel{\mathop{:}}n}\bigr{)}}{\partial q_{s}}$
@@ 6.1 Experimental Setup
    COMPAS dataset with 12 features and a binary label on whether a subject has recidivism in two years, where the sensitive attribute is the binary race feature [^1] https://github.com/propublica/compas-analysis. To simulate a setting with imbalanced sensitive attribute distribution, we considered 2500 training and 750 test samples, in both of which 80% are from $Z=0$non-Caucasian" and 20% of the samples are from $Z=1$Caucasian".
@@
    CelebA Proposed by [27], containing the pictures of celebrities with 40 attribute annotations, where we considered "gender" as a binary label, and the sensitive attribute is the binary variable on blond/non-blond hair. In the experiments, we used 5k training samples and 2k test samples. To simulate an imbalanced sensitive attribute distribution, 80% of both training and test samples are marked with Blond hair and 20% samples are marked with non-blond hair.
DP-based Learning Methods: We performed the experiments using the following DP-based fair classification methods: 1) DDP-based KDE method [6] and FACL [12], 2) the mutual information-based fair classifier [11], 3) the maximal Correlation-based RFI classifier [13], to learn binary classification models on COMPAS and Adult datasets. For CelebA experiments, we used the following two DP-based fair classification methods: KDE method [6], and mutual information (MI) fair classifier [11].
In the experiments, we attempted both a logistic regression classifier with a linear prediction model and a neural net classifier.
The neural net architecture was 1) for the COMPAS case, a multi-layer perceptron (MLP) with 4 hidden layers with 64 neurons per layer, 2) for the Adult case, an MLP with 4 hidden layers with 512 neurons per layer, 3) for the CelebA case, the ResNet-18 [28] architecture suited for the image input in the experiments.
Evaluation criteria: To evaluate the trained models, we used the averaged accuracy rate (Acc) as the classification performance metric and the Difference of Demographic Parity (DDP) as the fairness metric. Moreover, to quantify the bias effects of fair learners, we measured the negative rate (NR) conditioned to a sensitive attribute defined as $\mathrm{NR}(s)\mathrel{\mathop{:}}=P\bigl{(}\hat{Y}=0\mid S=s\bigr{)}$. This metric is defined to quantify the variations in prediction outcomes across subgroups with different sensitive attribute values.
@@ 6.2 Inductive Biases of Models trained in DP-based Fair Learning
Figure 3: The results of Figure 3’s experiments for a linear logistic regression model.
As the evaluated accuracy and DDP values in Figures 3,3 indicate, the DP-based fair learning algorithms managed to significantly reduce the DDP fairness violation while compromising less than 2% in accuracy. On the other hand, the negative rate ($\widehat{Y}=0$ prediction rate) across the two outcomes $S=0,\,S=1$ of the sensitive attribute tend toward the majority sensitive attribute as the DP-based fairness regularization became stronger, suggesting the conditional distribution of the prediction $\widehat{Y}$ given different sensitive attribute outcome $Z=z$’s moved closer to that of the majority sensitive attribute. The observed behavior held similarly using both the linear logistic regression model in Figure 3 and the neural net classifiers in Figure 3. Furthermore, this observation can also be shown on CelebA dataset using the ResNet-based model in Figure 6. The visualized image classification results are provided in Figure 6, and both figures for CelebA dataset can be viewed in the Appendix.
DRO-based Fair Learning. We tested Algorithm 1 utilizing a sensitive attribute-based distributional robust optimization (SA-DRO) to DP-based fair learning algorithms. In our experiments, we applied the SA-DRO algorithm to the DDP-based KDE fair learning algorithm proposed by [11], and RFI proposed by [13]. We kept the fairness regularization penalty coefficient to be $\lambda=0.9$. The DRO regularization coefficient can take over the range $\left[0,1\right]$, in this table, we set $\epsilon=0.9$ for SA-DRO case. In these figures 4, we applied the SA-DRO algorithm to the DDP-based KDE fair learning algorithm by [11], and RFI by [13]. We kept the fairness regularization penalty coefficient to be $\lambda=0.9$. The DRO regularization coefficient can take over the range $\left[0,1\right]$.
As Table 1 shows, we observed that the proposed SA-DRO reduces the tendency of the fair learning algorithm toward the majority sensitive attribute, and the resulting negative prediction rates conditioned to sensitive attribute outcomes became closer to the midpoint between the majority and minority conditional accuracies. On the other hand, the SA-DRO-based algorithms still achieve a low DDP value while the accuracy drop is less than 1%. Moreover, we visualize the prediction shifting in Figure 4 by applying the SA-DRO algorithm to the DDP-based KDE fair learning algorithm by [11], and RFI by [13]. We kept the fairness regularization penalty coefficient to be $\lambda=0.9$. The DRO regularization coefficient takes over the range $\left[0,1\right]$.
@@
Client 1 (Minority) Client 2-4 (Majority) Acc($\uparrow$) DDP($\downarrow$) Acc($\uparrow$) DDP($\downarrow$) FedAvg 82.5% 0.208 90.3% 0.206 ERM(Local) 81.6% 0.203 89.0% 0.246 FedKDE 74.8% 0.022 89.9% 0.029 FedFACL 74.5% 0.014 89.7% 0.031 SA-DRO-KDE 79.3% 0.041 89.6% 0.042 SA-DRO-FACL 79.0% 0.055 89.2% 0.038 KDE(Local) 79.0% 0.032 88.2% 0.014 FACL(Local) 79.1% 0.025 88.6% 0.017
Client 1 (Minority) Client 2-5 (Majority)
@@
FedAvg 94.8% 0.380 94.3% 0.419
ERM(Local) 91.8% 0.374 90.3% 0.396
FedKDE 65.6% 0.088 88.8% 0.060
FedFACL 67.0% 0.068 88.5% 0.054
SA-DRO-KDE 69.0% 0.063 88.1% 0.062
SA-DRO-FACL 68.5% 0.057 87.7% 0.069
KDE(Local) 69.7% 0.055 87.6% 0.043
FACL(Local) 69.1% 0.067 87.5% 0.050
Table 2: Accuracy and DDP on Adult dataset
Table 3: Accuracy and DDP on CelebA dataset
Figure 4: Accuracy, DDP, and $\mathrm{NR}(s)$ shifting by applying SA-DRO on the COMPAS (upper) and Adult (lower) Datasets for both KDE and RFI. The first and the second column show that the accuracy and DDP among the whole groups or different subgroups are slightly affected. The third shows that after increasing the DRO regularization parameters, the weight of losses in majority and minority groups will converge to the original distribution. The last shows that the $\mathrm{NR}(s)$ for different subgroups will shift from the majority group to the midpoint between the minority group and the majority group to reduce the inductive biases.
For the baseline federated learning method with no fairness regularization, we utilized the FedAvg algorithm [29]. For the DP-based fair federated learning algorithms, we attempted the DDP-based KDE and FACL algorithms which result in single-level optimization problem and hence can be optimized in a distributed learning problem by averaging as in FedAvg. We refer to the extended federated learning version of these algorithms as FedKDE and FedFACL. We also tested our SA-DRO implementations of FedKDE and FedFACL, as well as the localized ERM, KDE, FACL models where each client trained a separate model only on her own data.
As our numerical results in Table 3 and Table 3 indicate, the inductive biases of DP-based federated learning could considerably lower the accuracy of Client 1 with a different majority sensitive attribute compared to the other clients. The accuracy drop led to a lower accuracy compared to Client 1’s locally fair trained model without any collaboration with the other clients, which may affect the client’s incentive to participate in the federated learning process. On the other hand, the SA-DRO implementations of the KDE and FACL methods achieved a better accuracy than Client 1’s local model while preserving the accuracy for the majority clients.

### v2

@@ 5 A Distributionally Robust Optimization Approach to DP-based Fair Learning
1:  Input: Training data $\{(\mathbf{x}_{i},y_{i},s_{i})_{i=1}^{n}\}$, parameters $\lambda,\delta\geq 0$, divergence $d$, dependence measure $\rho$, stepsizes $\alpha_{w},\alpha_{q}>0$, running iterations $T>0$
@@
${g_{\mathbf{q}}}_{s}=\frac{1}{n}\sum_{i:s_{i}=s}\bigl{[}\ell\bigl{(}f_{\mathbf{w}}(\mathbf{x}_{i}),y_{i}\bigr{)}\bigr{]}+\lambda\frac{\partial\rho\bigl{(}f_{\mathbf{w}}(\mathbf{x}_{1:n}),s_{1:n}\bigr{)}}{\partial q_{s}}$
@@
$\min_{f\in\mathcal{F}}\max_{Q_{S}:d(Q_{S},P_{S})\leq\delta}\mathbb{E}_{P_{\mathbf{X},Y|S}\cdot Q_{S}}\Bigl{[}\ell_{0/1}\bigl{(}\widehat{Y},Y\bigr{)}\Bigr{]}+\lambda\rho\bigl{(}\widehat{Y},S\bigr{)}$ (4)
According to this formulation, we solve the Lagrangian version of optimization problem (1) when $S$’s marginal distribution $q_{S}$ leads to the worst-case fair-regularized risk function in a distribution ball $\bigl{\{}q_{S}:\,d(q_{S},p_{S})\leq\delta\bigr{\}}$ where $d$ is a distance measure between probability distributions. In this formulation, we consider assigning different weights to samples with different sensitive attributes, which may result in different majority sensitive attributes. Since we are optimizing the worst-case performance over the distribution ball with a $\delta$ radius, the inductive biases discussed in the previous would become less effective under a greater $\delta$.
The proposed SA-DRO formulation results in Algorithm 1 which applies projected gradient descent ascent (GDA) to solve the minimax optimization problem in (4). Here, we use a parameterized classifier $f_{\mathbf{w}}$ to apply a gradient-based training algorithm. Also, the distance $d$ can be chosen as any standard $f$-divergence. In our experiments, we attempted the $\chi^{2}$-divergence divergence, which has been well-explored in the literature [Namkoong and Duchi, 2016, Bertsimas et al., 2019, Rahimian and Mehrotra, 2019]. Furthermore, a Lagrangian form of the SA-DRO problem (4) can be considered where the DRO constraint on $Q_{S}$ is transferred to the inner maximization objective function as $-\zeta d(Q_{S},P_{S})$ for a Lagrangian coefficient $\zeta>0$.
@@ 6.1 Experimental Setup
    COMPAS dataset with 12 features and a binary label on whether a subject has recidivism in two years, where the sensitive attribute is the binary race feature111 https://github.com/propublica/compas-analysis. To simulate a setting with imbalanced sensitive attribute distribution, we considered 2500 training and 750 test samples, in both of which 80% are from $S=0$non-Caucasian" and 20% of the samples are from $S=1$Caucasian".
@@
    CelebA Proposed by [Liu et al., 2018], containing the pictures of celebrities with 40 attribute annotations, where we considered "gender" as a binary label, and the sensitive attribute is the binary variable on blond/non-blond hair. In the experiments, we used 5k training samples and 2k test samples. To simulate an imbalanced sensitive attribute distribution, 80% of both training and test samples are marked with Blond hair and 20% samples are marked with non-blond hair.
DP-based Learning Methods: We performed the experiments using the following DP-based fair classification methods: 1) DDP-based KDE method [Cho et al., 2020a] and FACL [Mary et al., 2019], 2) the mutual information-based fair classifier [Cho et al., 2020b], 3) the maximal Correlation-based RFI classifier [Baharlouei et al., 2019], to learn binary classification models on COMPAS and Adult datasets. For CelebA experiments, we used the following two DP-based fair classification methods: KDE method [Cho et al., 2020a], and mutual information (MI) fair classifier [Cho et al., 2020b].
In the experiments, we attempted both a logistic regression classifier with a linear prediction model and a neural net classifier. The neural net architecture was 1) for the COMPAS case, a multi-layer perceptron (MLP) with 2 hidden layers with 128 neurons per layer, 2) for the Adult case, an MLP with 4 hidden layers with 512 neurons per layer, 3) for the CelebA case, the ResNet-18 [He et al., 2016] architecture suited for the image input in the experiments.
Evaluation criteria: To evaluate the trained models, we used the averaged accuracy rate (Acc) as the classification performance metric and the Difference of Demographic Parity (DDP) as the fairness metric. Moreover, to quantify the bias effects of fair learners, we measured the negative rate (NR) conditioned to a sensitive attribute defined as $\mathrm{NR}(s):=P\bigl{(}\hat{Y}=0\mid S=s\bigr{)}$. This metric is defined to quantify the variations in prediction outcomes across subgroups with different sensitive attribute values.
@@ 6.2 Inductive Biases of Models trained in DP-based Fair Learning
Figure 2: The first two columns show the trade-off between accuracy and DDP on the COMPAS and Adult dataset by applying LR-based fair classification methods, while the third column shows that the $\mathrm{NR}(s)$ for each subgroup $s\in\{0,1\}$ will converge to near the majority sensitive attribute.
Figure 3: Both (a) and (b) show the trade-off between accuracy and DDP on the imbalanced CelebA dataset by applying MI fair classification method, while (c) shows that the $\mathrm{NR}(s)$ for each subgroup will converge to the majority, thus causing more discrimination on the minority group.
(a) ERM Classifier for CelebA
(b) Fair Classifier for CelebA
Figure 5: Accuracy, DDP, and $\mathrm{NR}(s)$ values attained by SA-DRO while varying the Lagrangian coefficient of the DRO regularization term on COMPAS (upper) and Adult (lower) datasets.
As the evaluated accuracy and DDP values in Figures 1 indicate, the DP-based fair learning algorithms managed to significantly reduce the DDP fairness violation while compromising less than 2% in accuracy. On the other hand, the negative rate ($\widehat{Y}=0$ prediction rate) across the two outcomes $S=0,\,S=1$ of the sensitive attribute tend toward the majority sensitive attribute as the DP-based fairness regularization became stronger, suggesting the conditional distribution of the prediction $\widehat{Y}$ given different sensitive attribute outcome $S=s$’s moved closer to that of the majority sensitive attribute. The observed behavior held similarly using both the linear logistic regression model in Figure 2 and Figure 3, and Figure 4 shows how the inductive biases lead to misclassification on CelebA dataset.
DRO-based Fair Learning. We tested Algorithm 1 utilizing a sensitive attribute-based distributional robust optimization (SA-DRO) to DP-based fair learning algorithms. In our experiments, we applied the SA-DRO algorithm to the DDP-based KDE fair learning algorithm proposed by Cho et al. [2020b], and RFI proposed by Baharlouei et al. [2019]. We kept the fairness regularization penalty coefficient to be $\lambda=0.9$. Following the commonly-used implementation of DRO, we used a Lagrangian penalty term $-\zeta d(P_{S},Q_{S})$ in the inner maximization problem to perform DRO. Therefore, the DRO regularization coefficient, also the Lagrangian multiplier $\zeta$, can take over the range $\left[0,+\infty\right]$, in the table 1, we set $\zeta=0.9$ for SA-DRO case. The visualized results for various DRO regularization coefficients can be found in Appendix.
@@
As Table 1 shows, we observed that the proposed SA-DRO reduces the tendency of the fair learning algorithm toward the majority sensitive attribute, and the resulting negative prediction rates conditioned to sensitive attribute outcomes became closer to the midpoint between the majority and minority conditional accuracies. On the other hand, the SA-DRO-based algorithms still achieve a low DDP value while the accuracy drop is less than 1%.
In Figure 5, we visualized the results of applying SA-DRO algorithm to the DP-based KDE by Cho et al. [2020b], and RFI by Baharlouei et al. [2019] for various DRO coefficients. We kept the fairness regularization penalty coefficient to be $\lambda=0.9$, and the DRO regularization coefficient took over the range $\left[0,1\right]$. This Figure 5 shows that the accuracy and DDP among the whole groups or different subgroups are slightly affected, while the $\mathrm{NR}(s)$ for different subgroups will shift from the majority group to the midpoint between the minority group and the majority group to effectively reduce the inductive biases.
For the baseline federated learning method with no fairness regularization, we utilized the FedAvg algorithm [McMahan et al., 2017]. For the DP-based fair federated learning algorithms, we attempted the DDP-based KDE and FACL algorithms which result in single-level optimization problem and hence can be optimized in a distributed learning problem by averaging as in FedAvg. We refer to the extended federated learning version of these algorithms as FedKDE and FedFACL. We also tested our SA-DRO implementations of FedKDE and FedFACL, as well as the localized ERM, KDE, FACL models where each client trained a separate model only on her own data.
To show the impacts of such inductive biases in practice, we focused on a setting with heterogeneous sensitive attribute distributions across clients where the clients’ majority sensitive attribute outcome may not agree. Figure 6 illustrates such a federated learning scenario over the Adult dataset, where Client 1’s majority sensitive attribute (female samples) is different from the network’s majority group (male samples). In the experiment, Client 1’s test accuracy with a DP-based fair federated learning was significantly lower than the test accuracy of a locally-trained fair model learned only on Client 1’s data. Such numerical results suggest the possibility of the minority clients’ lack of incentive to participate in the fair federated learning process.
To test our proposed DRO approach, we applied the SA-DRO method. As our numerical results in Table 2 indicate, the inductive biases of DP-based federated learning could considerably lower the accuracy of Client 1 with a different majority sensitive attribute compared to the other clients. The accuracy drop led to a lower accuracy compared to Client 1’s locally fair trained model without any collaboration with the other clients, which may affect the client’s incentive to participate in the federated learning process. On the other hand, the SA-DRO implementations of the KDE and FACL methods achieved a better accuracy than Client 1’s local model while preserving the accuracy for the majority clients and maintaining the same level of DDP no more than 0.05. We found similar results in the CelebA federated learning experiments, as in Table 3 in the Appendix.
@@ 6.3 DP-based Fair Classification in Heterogeneous Federated Learning
Client 1 (Minority) Client 2-4 (Majority)
@@
FedAvg 82.5% 0.208 90.3% 0.206
ERM(Local) 81.6% 0.203 89.0% 0.246
FedKDE 74.8% 0.022 89.9% 0.029
FedFACL 74.5% 0.014 89.7% 0.031
SA-DRO-FedKDE 79.3% 0.041 89.6% 0.042
SA-DRO-FedFACL 79.0% 0.049 89.1% 0.036
KDE(Local) 79.0% 0.032 88.2% 0.014
FACL(Local) 79.1% 0.025 88.6% 0.017